OpenStudy (anonymous):

What is the simplified form of 8x / x^2+4x+4 divided by 4x^2 /x^2-4 ?

jhonyy9 (jhonyy9):

can you factorizing ?

OpenStudy (anonymous):

When dividing by rational functions, you should actually multiply by the reciprocal(flip it).

OpenStudy (anonymous):

it is (x-2)/2x(x+2) correct?

OpenStudy (anonymous):

\[[8x \div(x ^{2} +4x+4) ] \times[(x ^{2}-4)/4x ^{2}]\]

jhonyy9 (jhonyy9):

sashley - so yes but for you can simplifie you need factoriz it completly - if this is possibile

OpenStudy (anonymous):

yeah

jhonyy9 (jhonyy9):

x^2 +4x +4 = ? x^2 -4 = ?

jhonyy9 (jhonyy9):

and you need to know again how you divide two fractions

OpenStudy (anonymous):

x^2 +4x +4 = (x+2)(x+2) Difference of squares x^2 -4 = (x-2)(x+2)

jhonyy9 (jhonyy9):

sashley -right and the next step ?

jhonyy9 (jhonyy9):

how you divide two fractions ?

OpenStudy (anonymous):

look at my posts above

jhonyy9 (jhonyy9):

for ornadios can you rewrite it step by step ? please